Your firm currently has 1,602,462 shares of stock outstanding that sell for 35.00. Additionally, you've issued 14,293 bonds, with a par value of S1,000. They currently sell for 96.8% of par.

What is the firm's weight of debt?

(Enter your answer as a percentage to 1 decimal place. ex: 12.3)
Answer #1

Value of Equity = Number of Shares(Stock Price)

Value of equity = 1,602,462(35) = $56,086,170

Value of Debt = Number of Bond(Bond Price)

Value of Debt = (14,293)(1,000*0.968) = $13,835,624

Weight of Debt = (13,835,624)/(13,835,624 + 56,086,170)

Weight of Debt = 0.198

Weight of Debt = 19.8%